OpenStudy (anonymous):

If the endpoints of the diameter of a circle are 3,10 and -1,6 what are the coordinates of its center?

OpenStudy (deoxna):

Find the midpoint of that line, You can do this by finding the average of each coordinate, so (3+(-1))/2 = 1 (10+6)/2 = 8 So the center is at (1,8)